OpenStudy (mandii):

What are the solutions to the following system of equations? y = x2 + 12x + 35 4x + y = –28

jhonyy9 (jhonyy9):

y=-28-4x y=x2+12x+35 28-4x=0 28=4x x=28/4 =7 x=7 x2+12x+35=0 x_1,_2=(-12+/-sqrt(144-140))/2=(12+/-sqrt4)/2=5 and 7